(1 point) Find the least-squares solution x* of the system [10] 0 1 x= 4 Loo] [-9(1 point) Find the least-squares solution x* of the system [ 1 | -1 ( 5 -2] [ 10 ] 2 X= 2 . 4 ] (-36]
Answer #1

(1) We have given:

\begin{bmatrix} 1 & 0\\ 0 & 1\\ 0 & 0 \end{bmatrix}x = \begin{bmatrix} 6 \\ 4 \\ -9 \end{bmatrix} Which can be written as: Ax = b where, A = \begin{bmatrix} 1 & 0\\ 0 & 1\\ 0 & 0 \end{bmatrix} and b = \begin{bmatrix} 6 \\ 4 \\ -9 \end{bmatrix} .

Multiplying both side of Ax = b by A^T we get:

A^TAx = A^Tb

\implies \begin{bmatrix} 1 & 0 & 0\\ 0 & 1 & 0 \end{bmatrix}\begin{bmatrix} 1 & 0\\ 0 & 1\\ 0 & 0 \end{bmatrix}x = \begin{bmatrix} 1 & 0 & 0\\ 0 & 1 & 0 \end{bmatrix}\begin{bmatrix} 6\\ 4\\ -9 \end{bmatrix}

\begin{bmatrix} 1 & 0\\ 0 & 1 \end{bmatrix} x = \begin{bmatrix} 6\\ 4 \end{bmatrix}

\implies x^* = \begin{bmatrix} 6\\ 4 \end{bmatrix}.

(2)

We have given \begin{bmatrix} 1 & -2\\ -1 & 2\\ 5 & 4 \end{bmatrix}x = \begin{bmatrix} 10\\ 2\\ -36 \end{bmatrix} which can be written as Ax = b where A = \begin{bmatrix} 1 & -2\\ -1 & 2\\ 5 & 4 \end{bmatrix} and b = = \begin{bmatrix} 10\\ 2\\ -36 \end{bmatrix} .

Multiplying both side of Ax = b by A^T we get:

A^TAx = A^Tb

\implies \begin{bmatrix} 1 & -1 & 5\\ -2 & 2 & 4 \end{bmatrix}\begin{bmatrix} 1 & -2\\ -1 & 2\\ 5 & 4 \end{bmatrix}x = \begin{bmatrix} 1 & -1 & 5\\ -2 & 2 & 4 \end{bmatrix} \begin{bmatrix} 10\\ 2\\ -36 \end{bmatrix}

\begin{bmatrix} 27 & 16\\ 16 & 24 \end{bmatrix}x = \begin{bmatrix} -172\\ -160 \end{bmatrix} solving this system of two equation in two unknows we get:

x^* = \begin{bmatrix} -4\\ -4 \end{bmatrix}.